# ConVerSum: A Contrastive Learning based Approach for Data-Scarce Solution of Cross-Lingual Summarization Beyond Direct Equivalents

### v1

@@ Abstract
Cross-lingual summarization (CLS) is a sophisticated branch in Natural Language Processing that demands models to accurately translate and summarize articles from different source languages. Despite the improvement of the subsequent studies, This area still needs data-efficient solutions along with effective training methodologies. To the best of our knowledge, there is no feasible solution for CLS when there is no available high-quality CLS data. In this paper, we propose a novel data-efficient approach, ConVerSum, for CLS leveraging the power of contrastive learning, generating versatile candidate summaries in different languages based on the given source document and contrasting these summaries with reference summaries concerning the given documents. After that, we train the model with a contrastive ranking loss. Then, we rigorously evaluate the proposed approach against current methodologies and compare it to powerful Large Language Models (LLMs)- Gemini, GPT 3.5, and GPT 4 proving our model performs better for low-resource languages’ CLS. These findings represent a substantial improvement in the area, opening the door to more efficient and accurate cross-lingual summarizing techniques.
Keywords Cross-lingual  $\cdot$ summarization  $\cdot$ contrastive learning  $\cdot$ candidate summary  $\cdot$ reference summary  $\cdot$ low-resource
1 Introduction
@@
Cross-lingual Summarization (CLS) is the process of producing a summary in a different language from the source language. This can aid in easily acquiring the gist of a foreign language article. This task can be thought of as a fusion of machine translation (MT) and monolingual summarization (MS), two open natural language processing (NLP) challenges that have been researched for decades [1] [2]. CLS is an extremely complex process from the perspective of both data and models. In terms of data, as opposed to MS, the lack of naturally occurring documents in a source language and comparable summaries in several target languages make it difficult to gather large-scale and human-annotated datasets [3] [4]. From the viewpoint of models, CLS necessitates the ability to translate and summarize, making it challenging to provide precise summaries using CLS independently [5]. Although the absence of high-quality large-scale data has recently been addressed [6], the CLS still needs models that take data-efficient solutions into account.
@@
Previous attempts at CLS used pipeline methods such as translate-then-summarize [7] and summarize-then-translate [8]. Due to the necessity for several models, these methods are not only computationally costly but also experience error propagation [9] among models, which diminishes overall performance rendering it inappropriate for the scenario in real life [3]. To tackle this problem, CLS has recently been accomplished by utilizing Sequence-to-Sequence(Seq2Seq) [10] [11] and Transformers-based models [12] to perform the summarization in a single model [9] [13] [6]. The aforementioned achievements have paved the way for the field of CLS research and progressively fueled curiosity in CLS. However, no one proposed any CLS solution when there is no available large-scale CLS corpus to the best of our knowledge. Thus, we wish to investigate the question: Can we develop a CLS model which will be able to generate coherent and informative summaries when no CLS data is available?
@@
In this paper, we propose a model ConVerSum which provides a data-scarce solution for CLS without parallel corpora by leveraging contrastive learning. At first, we generate versatile candidate summaries in different languages using a seq2seq model, use diverse Beam search to explore word sequences, measure the quality of candidate summaries, apply contrastive learning to minimize positive and negative pairs and train the model with a contrastive ranking loss to improve similarity scores. ConVerSum is evaluated using standard metrics and tested on different cross-lingual language pairs and LLMs. Therefore, our main contribution to this paper is as follows:
@@
  • We carry out a comprehensive comparative study to identify the best-performing approach and prove ConVerSum’s robustness on different cross-lingual pairs and low-resource languages. We evaluate it with state-of-the-art LLMs like GPT 3.5, GPT 4, and Gemini.
2 Related Works
Early CLS work utilized pipeline approaches, dividing CLS into MS and MT sub-tasks and completing them sequentially. These approaches can be classified as summarize- then-translate [14, 8] or translate-then-summarize [7, 15, 16, 17] based on the sequence of completed sub-tasks. End-to-end methods that performed CLS with a single model achieve popularity with the emergence of neural models. These methods can be categorized into four frameworks [18] such as multi-task framewok [9, 5, 19], knowledge-distillation framework [20, 21], resource-enhanced framework [22, 23, 22] and pre-training framework [24, 25].
Contrastive learning has recently successfully emerged in various fields of NLP [26]. The application of contrastive learning for abstractive summarization is a nascent yet promising branch of research [27, 28, 29].
3 ConVerSum
In this section, we elaborate on our proposed approach ConVerSum. At first, candidate summaries are generated in different languages using monolingual languages. We generate candidate summaries inspired by the candidate generation process of SimCLS [27]. From a monolingual document, numerous candidate summaries in several languages are generated using the seq2seq model, mT5 [25]. Because it was extensively trained on a diverse language dataset, it possesses multilingual skills. The model can be easily adjusted to different language contexts. This scenario makes use of the improved many-to-many (m2m) mT5 checkpoint111csebuetnlp/mT5_m2m_crossSum_enhanced. However not all summaries that are generated are of excellent quality, and it’s possible for repeating summaries to be produced. This is addressed using diverse Beam search to generate candidate summaries, exploring various word sequences and producing high-quality summaries.
Then, the generated candidate summaries are evaluated concerning the reference summaries and the given documents.The evaluation function for candidate summaries is initiated with XLM-RoBERTa [30] base model [^2] FacebookAI/xlm-roberta-base for mapping semantically similar text from several languages into a shared embedding space. This procedure captures the meaning and context of the text. Cosine similarity is used to determine the similarity of candidate summaries and reference summaries to the supplied source document. LaSE [6], a scoring method devised exclusively for assessing CLS, is calculated for each candidate summary and ranked in descending order. As a result, the highest-scoring candidate is more similar to the reference summary, while the lowest-scoring candidate is less comparable. Both cosine similarity and LaSE scores are employed to capture many aspects of summary quality, ensuring a full evaluation as well as the correctness and relevancy of the summaries. This is crucial for tasks requiring accurate factual information and contextual knowledge.
Next, we implement contrastive learning objective for our task. The goal of the contrastive learning is to minimize the distance between positive pairs and maximize the distance between negative pairings in order to improve cross-lingual summary generation. Using a contrastive ranking loss function, the model is trained to push negative pairings further away and bring positive pairs closer together. By performing this, it is ensured that the summaries that are produced are relevant to the context and correctly semantically. To achieve this goal, we construct positive and negative pairs. Positive pairings reflect summaries that are more similar to the reference summary concerning documents, implying greater quality. In contrast, negative pairings comprise summaries that are of inferior quality when compared to other candidate summaries and the reference summary. This method aids in discriminating between superior and worse summaries, increasing the model’s ability to generate high-quality summaries.
The model is trained on cross-lingual summary generation with contrastive ranking loss to capture complex connections between the original documents and their related summaries in different languages. It learns to assign higher similarity scores to positive summary pairings and lower scores to negative combinations. The model’s efficacy is assessed using the LaSE and BERTScore. The model’s resilience is tested on the existing CLS dataset and low-resource monolingual datasets, and performance is compared to the most powerful current LLMs.
Figure 1 illustrates the general structure of ConVerSum. Furthermore, we repeat the whole process using Flan-T5 [31] instead of mT5 where we use the checkpoint333google/flan-t5-large to identify which approach performs better. All the details are explained in Section 4.
@@ 3 ConVerSum
Figure 1: General Structure of ConVerSum. Here CS, RS, Lang and Doc refer to candidate summary, reference summary, language and document, respectively. For better realization, we assume score1 > score2 > score3 > score4.
4 Experimental Evaluation
4.1 Training Setup

### v2

# ConVerSum: A Contrastive Learning-based Approach for Data-Scarce Solution of Cross-Lingual Summarization Beyond Direct Equivalents

@@ Abstract.
Cross-lingual summarization (CLS) is a sophisticated branch in Natural Language Processing that demands models to accurately translate and summarize articles from different source languages. Despite the improvement of the subsequent studies, This area still needs data-efficient solutions along with effective training methodologies. To the best of our knowledge, there is no feasible solution for CLS when there is no available high-quality CLS data. In this paper, we propose a novel data-efficient approach, ConVerSum, for CLS leveraging the power of contrastive learning, generating versatile candidate summaries in different languages based on the given source document and contrasting these summaries with reference summaries concerning the given documents. After that, we train the model with a contrastive ranking loss. Then, we rigorously evaluate the proposed approach against current methodologies and compare it to powerful Large Language Models (LLMs)- Gemini, GPT 3.5, and GPT 4o proving our model performs better for low-resource languages’ CLS. These findings represent a substantial improvement in the area, opening the door to more efficient and accurate cross-lingual summarizing techniques.
cross-lingual, summarization, low-resource, LLM, contrastive learning
copyright: noneconference: ; journalyear: ;doi: isbn:
1. Introduction
@@
Cross-lingual Summarization (CLS) is the process of producing a summary in a different target language (e.g. Bengali) from the source language (e.g. Japanese) (Figure 1). This can aid in easily acquiring the gist of a foreign language article. This task can be thought of as a fusion of machine translation (MT) and monolingual summarization (MS), two open natural language processing (NLP) challenges that have been researched for decades (Paice, 1990; Brown et al., 1993). CLS is an extremely complex process from the perspective of both data and models. In terms of data, as opposed to MS, the lack of naturally occurring documents in a source language and comparable summaries in several target languages make it difficult to gather large-scale and human-annotated datasets (Ladhak et al., 2020; Perez-Beltrachini and Lapata, 2021). From the viewpoint of models, CLS necessitates the ability to translate and summarize, making it challenging to provide precise summaries using CLS independently (Cao et al., 2020a). Although the absence of high-quality large-scale data has recently been addressed (Bhattacharjee et al., 2023), the CLS still needs models that take data-efficient solutions into account.
Figure 1. Example of CLS
[…]
@@
Previous attempts at CLS used pipeline methods such as translate-then-summarize (Leuski et al., 2003) and summarize-then-translate (Wan et al., 2010). Due to the necessity for several models, these methods are not only computationally costly but also experience error propagation (Zhu et al., 2019) among models, which diminishes overall performance rendering it inappropriate for the scenario in real life (Ladhak et al., 2020). To tackle this problem, CLS has recently been accomplished by utilizing Sequence-to-Sequence(Seq2Seq) (Cho et al., 2014; Sutskever et al., 2014) and Transformers-based models (Vaswani et al., 2017) to perform the summarization in a single model (Zhu et al., 2019; Cao et al., 2020b; Bhattacharjee et al., 2023). The aforementioned achievements have paved the way for the field of CLS research and progressively fueled curiosity in CLS. However, no one proposed any CLS solution when there is no available large-scale CLS corpus to the best of our knowledge. Thus, we wish to investigate the question: Can we develop a CLS model which will be able to generate coherent and informative summaries when no CLS data is available?
@@
In this paper, we propose a model ConVerSum which provides a data-scarce solution for CLS without parallel corpora by leveraging contrastive learning. At first, we generate versatile candidate summaries in different languages using a seq2seq model, use diverse Beam search to explore word sequences, measure the quality of candidate summaries, apply contrastive learning to minimize positive and maximize negative pairs and train the model with a contrastive ranking loss to improve similarity scores. ConVerSum is evaluated using standard metrics and tested on different cross-lingual language pairs and Large Language Models (LLMs). Therefore, our main contribution to this paper is as follows:
@@
  • We carry out a comprehensive comparative study to identify the best-performing approach and prove ConVerSum’s robustness on different cross-lingual pairs and low-resource languages. We evaluate it with state-of-the-art LLMs like GPT 3.5, GPT 4o, and Gemini.
2. Related Works
Over the past five years, more than twenty articles have been published by researchers who did their best to solve the CLS task (Wang et al., 2022). Moreover, contrastive learning in abstractive summarization has attracted the attention of researchers lately (Zhang et al., 2022). We analyze existing publications to identify the most sophisticated strategies employed in recent studies. This section provides a full discussion of relevant articles and their working techniques on CLS and contrastive learning in summarization.
2.1. Pipeline Methods
Early CLS work focused on pipeline approaches, which decompose CLS into MS and MT sub-tasks and complete them sequentially. These approaches can be classified as summarize-then-translate (Orăsan and Chiorean, 2008; Wan et al., 2010) or translate-then-summarize (Leuski et al., 2003; Wan, 2011; Boudin et al., 2011; Yao et al., 2015) based on the sequence of completed sub-tasks. The pipeline method is straightforward however has some drawbacks. First of all, they are computationally very expensive for using multiple models; secondly, these techniques are prone to error propagation, resulting in unsatisfactory overall performance (Zhu et al., 2019). Moreover, latency occurs during inference. Additionally, training MT models requires either a big corpus or commercial services.
2.2. End-to-End Methods
Advancements in seq2seq (Cho et al., 2014; Sutskever et al., 2014) and Transformer-based models (Vaswani et al., 2017) have led to the invention of end-to-end methods for performing CLS with a single model to resolve the drawbacks mentioned in Section 2.1. The end-to-end methods can be classified into four categories - multi-task framework, knowledge distillation framework, resource-enhanced framework and pre-training framework.
Multi-task is a machine learning approach in which a single model is designed by combining several related tasks. Several studies train unified models using CLS in conjunction with related tasks such as MT and MS (Zhu et al., 2019; Cao et al., 2020a; Takase and Okazaki, 2020; Bai et al., 2021; Liang et al., 2022). CLS models thus gain from the related tasks as well.
Knowledge distillation is a machine learning approach used to reduce a large, complicated model to a smaller, simpler one. This is accomplished by transferring information from the larger model, known as the teacher model, to the smaller model, known as the student model (Hinton et al., 2015). To teach the CLS model in the knowledge-distillation framework, several researchers employ either MT/MS or both models because to have a mutual connection. In addition to CLS labels, the student model can gain knowledge from the output or hidden state of the teacher’s models. Extensive MS and MT datasets are leveraged to train MS and MT models, respectively which are bidirectional GRU (Cho et al., 2014). These models act as teachers for a CLS student model, with KL-divergence aligning generation probabilities and maximizing the training objective (Ayana et al., 2018). Afterward, transformer (Vaswani et al., 2017) is used as the backbone for the CLS student model and the MS teacher model and further trains the student model (Duan et al., 2019a).
The resource-enhanced framework enriches input documents with extra resources. Output summaries are generated based on encoded and enriched information. Translation patterns in CLS are investigated by encoding source language texts using a transformer encoder (Zhu et al., 2020). The TextRank toolset (Mihalcea and Tarau, 2004) helps create article graphs by extracting important information from input sequences (Jiang et al., 2022). A transformer-based clue encoder and a graph neural network-based graph encoder are used, respectively, to encode these clues and article graphs. Ultimately, the translation distribution from (Zhu et al., 2020) is included in final summaries generated by a transformer decoder with dual cross-attention methods.
In the context of NLP, pre-trained models are neural network architectures that are extensively trained on textual data to identify linguistic patterns and semantic representations. These models are trained on extensive text corpora, which helps them understand word and phrase semantic meanings as well as contextual relationships. By using pre-trained NLP models, transfer learning can be accomplished more successfully. This is because the gained information can be fine-tuned on certain NLP tasks improving efficiency and performance. The pre-trained models revolutionize NLP (Qiu et al., 2020).
Recently, multi-lingual pre-trained generative models have performed exceptionally well in various multi-lingual NLP tasks. mBART (Liu et al., 2020) is a multilingual pre-trained model which is another variant of BART (Lewis et al., 2020) model developed by Facebook AI. mBART is pre-trained with denoising objectives based on the BART architecture and vast unlabeled multilingual datasets. Specifically, it is trained in over 25 languages text data. Initially excelling in MT (Liu et al., 2020), a recent study shows that it can outperform many multi-task CLS models on large-scale datasets with straightforward fine-tuning (Liang et al., 2022). Subsequently, the language processing capabilities of mBART are expanded from 25 to 50 languages with mBART-50 (Tang et al., 2021). Aside from the BART-style pre-trained models, mT5 (Xue et al., 2021) shines out as a multilingual version of T5 (Raffel et al., 2020), trained across 101 languages with the T5-style span corruption objective. Despite the remarkable performance of these general pre-trained models, they lack explicit cross-lingual supervision and instead rely only on denoising or span corruption objectives across several languages. They have thus far not fully explored their cross-lingual potential.
2.3. Contrastive Learning for Monolingual Abstractive Summarization
Contrastive learning has recently successfully emerged in various fields of NLP (Zhang et al., 2022), to learn good and discriminative representations. Contrastive learning is a self-supervised approach to training models to differentiate between samples that are similar versus dissimilar. This approach is popularized in the SimCLR framework (Chen et al., 2020), which uses contrastive loss to minimize the distance between representations for similar data points and maximize it for dissimilar points.
The application of contrastive learning for abstractive summarization is a nascent yet promising branch of research. The SimCLS framework is designed to bridge the gap between standard seq2seq models and learning objectives by utilizing contrastive learning to adapt to evaluation metrics with the learning objective (Liu and Liu, 2021). A contrastive attention mechanism is proposed for abstractive sentence summarization that builds on the seq2seq framework by combining conventional attention, which focuses on relevant parts of the source sentence, and opponent attention, which focuses on less relevant parts, with both trained in opposite directions using novel softmax and softmin functions (Duan et al., 2019b). COLO, a contrastive learning-based re-ranking framework for one-stage summarizing is introduced that aligns training and assessment by producing summaries based on summary-level scores without the requirement for extra modules or parameters (An et al., 2022). CALMS, a unified summarizing model is constructed which is trained with contrastive learning by using MLGSum, a large-scale multilingual summarizing corpus including 1.1 million articles and summaries in 12 languages (Wang et al., 2021). It incorporates contrastive sentence ranking and sentence-aligned replacement to improve extractive skills and align sentence-level representations across languages.
3. ConVerSum
In this section, we elaborate on our proposed approach ConVerSum which is categorized into several phases. Figure 2 illustrates the general structure of ConVerSum. ConVerSum deals with the data scarcity challenges where the existing models require high-quality large-scale datasets as discussed in Section 2.
[…]
@@ 3. ConVerSum
Figure 2. General Structure of ConVerSum. Here CS, RS, Lang and Doc refer to candidate summary, reference summary, language and document, respectively. For better realization, we assume score1 ¿ score2 ¿ score3 ¿ score4 …scoreN.
3.1. First Phase: Candidate Summary Generation in Different Languages
In this phase, candidate summaries are generated in different languages using monolingual languages. We generate candidate summaries inspired by the candidate generation process of SimCLS (Liu and Liu, 2021). Though SimCLS is developed for monolingual abstractive summarization, their candidate generation process motivates us to create a robust CLS model. The detailed procedure is as follows.
3.1.1. Seq2Seq Model Selection
A renowned seq2seq model, mT5 (Xue et al., 2021) is utilized to generate multiple candidate summaries in different languages from the given monolingual document. mT5 has multilingual capabilities as it is pre-trained on a vast and diversified multilingual dataset spanning more than 100 languages. This thorough training enables the model to create summaries in many languages from a single monolingual source, making it very versatile and adaptable to a variety of linguistic settings. The enhanced many-to-many(m2m) mT5 checkpoint111csebuetnlp/mT5_m2m_crossSum_enhanced fine-tuned on all cross-lingual pairs of CrossSum (Bhattacharjee et al., 2023) dataset is employed in this case.
3.1.2. Diverse Summary Exploration
Many candidate summaries can be generated by utilizing the seq2seq model as described in Section 3.1.1. However, not all generated candidate summaries are high-quality. Moreover, there is a possibility of generating repetitive summaries and all summaries in the same language. To tackle these issues, diverse Beam search (Vijayakumar et al., 2016) is used during the generation of candidate summaries which explores different sequences of words in the generated summary, contributes to generating diverse candidate summaries, and finds out the high-quality summaries.
3.2. Second Phase: Candidate Summary Quality Measurement
In this phase, the generated candidate summaries as described in Section 3.1 are measured concerning the reference summaries and the given documents. The detailed process is as follows:
3.2.1. Evaluation Function Initialization
The evaluation function is initiated using XLM-RoBERTa (Conneau et al., 2020). The base model of XLM-RoBERTa [^2] FacebookAI/xlm-roberta-base available in huggingface transformers library [^3] https://huggingface.co/ is specifically employed in this case. The purpose of utilizing XLM-RoBERTa is to map semantically similar text from various languages into a shared embedding space. Encoding the candidate summaries, reference summaries, and documents can capture the meaning and context of the text.
3.2.2. Similarity Calculation
For similarity calculation, cosine similarity is utilized (Rahutomo et al., 2012). The purpose of calculating similarity is to identify how similar the candidate summaries and the reference summary are to the given source document. The general representation of the similarities between the source document, reference summary, and candidate summaries is as follows where C, R and S represent the candidate summaries, reference summary, and the source document, respectively.
$\text{Similarity}=f(C,R,S)=\frac{C\cdot S+R\cdot S}{\|C\|\|S\|+\|R\|\|S\|}$
This score is utilized to measure the contrastive ranking loss described in Section 3.3.2.
3.2.3. Candidate Summary Scoring
While generating candidate summaries as described in Section 3.1, LaSE (Bhattacharjee et al., 2023), which is specifically designed for evaluating CLS is computed concerning the reference summary. That means a LaSE score is assigned to each candidate summary and sorted in descending order. This indicates the highest-scoring candidate is more similar to the reference summary and the lowest-scoring candidate is less similar to the reference summary. This process can be defined as follows where $C_{i}$ and R represent the i-th candidate summary and reference summary, respectively.
$\text{LaSE}(S_{1})\geq\text{LaSE}(S_{2})\geq\dots\geq\text{LaSE}(S_{n})=f(C_{i},R)\quad\text{for each}\ i\in\{1,2,\dots,n\}$
Both cosine similarity and LaSE scores are used in the evaluation of candidate summaries to capture various characteristics of summary quality. Every metric offers distinct perspectives that enhance one another, resulting in a more thorough assessment as there is a possibility of generating bias using only one scoring method as CLS is an extremely complex task. Striking a balance between semantic coherence and exact content overlap enables a thorough examination and guarantees that the summaries that are produced are correct and relevant. This thorough method produces summaries that are of higher quality and more dependability, which is important for tasks requiring accurate factual information in addition to contextual knowledge.
3.3. Third Phase: Contrastive Learning Approach
We apply a contrastive learning objective that cultivates a common representation space to improve the quality of cross-lingual summary generation. The objective of this approach is to maximize the distance between negative pairings and minimize the distance between positive pairs. The model is trained to push negative pairings farther apart and bring positive pairs closer together in the representation space by employing a contrastive ranking loss function. This ensures that the generated summaries are both semantically correct and pertinent to the given context. The model’s capacity to differentiate between and provide excellent summaries is much enhanced by this approach. In this section, the whole process is described elaborately.
3.3.1. Positive-Negative Pair Construction
To implement the contrastive learning objective, we need to construct the pairs that are used for positive and negative samples for contrastive learning. To properly evaluate the quality of candidate summaries, we use a positive-negative pair creation procedure. This entails computing the scores of candidate summaries for the provided documents. Positive pairings include candidate summaries that are more comparable to the reference summary, implying greater quality. In contrast, negative pairings comprise summaries that are of inferior quality when compared to other candidate summaries and the reference summary. This strategy aids in discriminating between better and inferior summaries, hence improving the model’s capacity to produce high-quality summaries.
3.3.2. Contrastive Ranking Loss
The goal of the Contrastive Ranking Loss approach is to maximize the correlations between rankings of various candidate summaries. This is accomplished by adding a margin, which acts as a threshold to control the amount of space between summaries in pairs. This margin makes sure that summaries that should be similar are represented by positive pairings, which are encouraged to have high similarity scores, while summaries that should be dissimilar are represented by negative pairs, which are encouraged to have low similarity scores. The approach efficiently separates pertinent and extraneous summaries by enforcing this margin. Furthermore, Contrastive Ranking Loss promotes proper ranking connections by punishing deviations from the intended ranking order. This means that summaries should be ordered based on their relevance or similarity to an expected reference point. By optimizing the loss function using these ranking connections, the technique guarantees that the most relevant summaries are prioritized in the output. This loss can be mathematically represented as follows:
$\mathcal{L}=\sum_{i=1}^{n}\max(0,\text{margin}-f(C_{i},R^{+})+f(C_{i},R^{-}))$
where, $R^{+}$ and $R^{-}$ represent the positive and negative pairs, respectively.
3.4. Fourth Phase: Model Training
The model is trained on the cross-lingual summary generation task by using the contrastive ranking loss described in Section 3.3.2. This process entails modifying the model’s parameters to successfully capture the complex connections between the source documents and their related summaries across multiple languages. This is accomplished by using Contrastive Loss as the optimization goal, which provides a margin to encourage the model to differentiate between important and irrelevant summary pairs. During training, the model learns to give high similarity scores to positive summary pairings that transmit comparable information across languages and low similarity scores to negative pairs that include different documents.
3.5. Fifth Phase: Fine-Tuning and Evaluation
By fine-tuning the model, it obtains the capacity to create short and informative summaries that keep the substance of the original material across language limitations, thereby increasing the performance of cross-lingual summarizing tasks. Then, the model’s effectiveness is evaluated by the standard evaluation metrics LaSE (Bhattacharjee et al., 2023) and BERTScore (Zhang* et al., 2020).
3.6. Sixth Phase: Model Verification
As we developed our model on a monolingual high-resource dataset, we need to verify the model’s robustness on the existing CLS dataset. Furthermore, we explore our approach to different low-resource monolingual datasets to ensure adaptability in scenarios when there is no parallel CLS data. Moreover, we make a performance comparison of our model to the most powerful latest LLMs.
Furthermore, we repeat the whole process using Flan-T5 (Chung et al., 2022) instead of mT5 where we use the checkpoint444google/flan-t5-large to identify which approach performs better. All the details are explained in Section 4.
4. Experimental Evaluation
4.1. Training Setup